A 1.2-in outer diameter pipe is to span across a river at 115ft-wide section while being completely immersed in water. The average flow velocity of the water is 8ft/s, and its temperature is 70°F. The drag coefficient is a) 0.85 b) 1.05 c)1.25 d) 0.95 e) 0.65
Answer #1

Please apprecate the work done by UPVOTING .. Please .. Thank you .

lbm At 70°cf Y = 6.556810-4 I ftsu B = 62-3 lbm 1413 This, By all giren things and above value, Reynolds No)= JUD M 62.3X 8 X